# Characterizing Evolution in Expectation-Maximization Estimates for Overspecified Mixed Linear Regression

### v1

@@ 1.1 Related Work
When ground truth parameters of some components in mixture models have no separation, this means we are using a model with more components to fit a ground truth data distribution with fewer components, which is called the overspecified setting. In particular, the 2GMM/2MLR model is considered overspecified when there is no separation between the true regression parameters or location parameters of its two components. Moreover, it has been observed that the convergence of EM can be prohibitively slow when dealing with poorly separated mixtures (Redner & Walker, 1984), while achieving rapid linear convergence or superlinear convergence in cases with strong separation (Kwon et al., 2021; Ghosh & Kannan, 2020). The convergence rate of regression parameters using EM for 2GMM in the overspecified setting, assuming known mixing weights, is analyzed in Dwivedi et al. (2020b); Dhawan et al. (2023). Kwon et al. (2019; 2021) derived iteration complexity and statistical accuracy for the known-weight case across all SNR regimes—but their results in the low-SNR setting apply only when mixing weights are known to be balanced. Kwon & Caramanis (2020) analyzed EM for MLR with more than two components under restrictive conditions—such as requiring initializations very near the ground truth and well-separated regression parameters—which do not cover cases of overspecification where the ground truth parameters coincide. Nevertheless, there is a lack of prior research formally analyzing the evolution of EM updates for 2MLR with unknown regression parameters and mixing weights in the overspecified setting, with most prior research primarily concentrating on scenarios with known mixing weights. The behavior of EM updates in overspecified settings is not fully understood when dealing with unknown mixing weights, whether they are balanced or unbalanced. In this paper, we address this gap by conducting a detailed analysis of the EM updates for overspecified 2MLR, based on the expressions of integrals with Bessel functions given in (Luo & Hashemi, 2024). Even though Luo & Hashemi (2024) provided closed-form EM update rules (using Bessel functions) for 2MLR across all SNR regimes, but they focus primarily on the high-SNR case and do not discuss convergence guarantees or the dynamics in the overspecified setting and the low-SNR regime. The statistical rates for parameter estimation of a class of overspecified finite mixture models have been studied in several studies (Chen, 1995; Ho & Nguyen, 2016; Heinrich & Kahn, 2018; Doss et al., 2020; Manole & Ho, 2020; 2022; Ho et al., 2022; Nguyen et al., 2023). The optimal minimax rates for 2MLR with known balanced mixing weights have been determined in Kwon et al. (2021), while those for 2GMM with known mixing weights have been established in Dwivedi et al. (2020b) using the scheme outlined by Balakrishnan et al. (2017). However, the statistical error for 2MLR with unknown mixing weights remains unclear. We develop new techniques to sharpen the analysis for this statistical error.
@@ 3 Motivating Examples
As motivating examples, we highlight the tasks of haplotype assembly in bioinformatics and genomics (Cai et al., 2016) and phase retrieval, which arises in numerous fields including acoustics, optics, and quantum information (Candes et al., 2015), to justify the practical implications of our work.
@@ 3.1 Haplotype Assembly
Haplotypes are sequences of chromosomal variations in an individual’s genome that are crucial for determining the individual’s disease susceptibility. Haplotype assembly involves reconstructing these sequences from a mixture of sequenced chromosome fragments. Notably, humans have two haplotypes, i.e., they are diploid organisms (see Cai et al. (2016) for a clear mathematical formulation of the problem). For diploids, the primary challenge is to reconstruct two distinct haplotypes (binary sequences of single nucleotide polymorphisms—SNPs) from short, noisy sequencing reads. Each read corresponds to a local window of the genome but originates from one of the two chromosomes. The ambiguity in the haplotype origin of each read can be modeled as a mixture of two linear regression models with symmetric parameters, aligning with the model discussed in our work. Following Cai et al. (2016), let $\theta^{\ast}\in\{-1,+1\}^{d}$ represent one haplotype, and the other haplotype is its negation $\theta^{\ast}$. The sequencing data comprises $n$ noisy reads $\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\{0,1\}^{d}$ is a sparse binary indicator vector encoding which SNPs are covered by the read, and $y_{i}\in\mathbb{R}$ is the observed read signal (e.g., aggregate allele value), modeled as
$y_{i}=(-1)^{z_{i}+1}\langle x_{i},\theta^{\ast}\rangle+\varepsilon_{i},$ (9)
where $z_{i}\in\{1,2\}$ is a latent variable indicating whether the read originates from haplotype $\theta^{\ast}$ or $\theta^{\ast}$, and $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ is Gaussian noise. In bioinformatics and genomics, sequencing data is inherently noisy and often exhibits a low SNR. This is due to short read lengths relative to haplotype length and high sequencing error rates in biological variation (e.g., substitution and phasing errors), both of which increase the effective noise variance $\sigma^{2}$. As noted in Sankararaman et al. (2020), the efficacy of EM-type methods is demonstrated for this problem—indeed, the authors remark that “this iterative update rule is reminiscent of the class of Expectation Maximization algorithms.” Our work aligns with the diploid haplotype assembly problem under high noise conditions, and for precise mathematical derivations, we leverage the Gaussianity assumption on $x_{i}$ to characterize EM behavior in the symmetric two-component mixture of linear regressions (2MLR) setting.
@@ 3.2 Phase Retrieval
$y^{\prime}_{i}=y_{i}^{2}-\sigma^{2}=|\langle x_{i},\theta^{\ast}\rangle|^{2}+\underbrace{\left[2(-1)^{z_{i}+1}\langle x_{i},\theta^{\ast}\rangle\varepsilon_{i}+(\varepsilon_{i}^{2}-\sigma^{2})\right]}_{\xi_{i}}=|\langle x_{i},\theta^{\ast}\rangle|^{2}+\xi_{i}.$ (10)
This formulation is essentially the phase retrieval model with a heteroskedastic error term $\xi_{i}$, which has zero mean, i.e., $\mathbb{E}[\xi_{i}]=0$. Therefore, by leveraging our results on symmetric 2MLR, we can directly establish convergence guarantees for the corresponding phase retrieval problem.
@@ 6 Finite-Sample Level Analysis
Proposition 6.4 (see proof in Appendix F, Subsection F.4) demonstrates our sample complexity $n=\Omega(d\vee\log\frac{1}{\delta})$, whereas Lemma 1 of Dwivedi et al. (2020b) presents a sample complexity $n=\Omega(d\log\frac{1}{\delta})$. The difference lies in the techniques used: their multiplicative log factor is derived through the standard symmetrization technique with Rademacher variables and the application of the Ledoux-Talagrand contraction inequality (see the proof of Lemma 1 on page 44 of Dwivedi et al. (2020b) and Appendix E of Kwon et al. (2021) for Lemma 11). In contrast, we obtain our additive log factor by applying the rotational invariance of Gaussians and expressing the $\ell_{2}$ norm of the orthogonal error as the geometric mean of two Chi-square distributions (see Appendix E, Appendix E.3).
@@ 8 Experiments
In this section, we validate our theoretical findings in the previous sections with numerical experiments.
@@ 9 Conclusions
Building upon the analysis and established connections between the diffusion model objective and the classic EM algorithm in GMM (Shah et al., 2023), we foresee extending this analysis from GMM to MLR and establishing the time and sample complexities involved in learning the diffusion model objective, as discussed in recent works (Chen et al., 2024a; Gatmiry et al., 2024). The practical applications (haplotype assembly (Cai et al., 2016; Sankararaman et al., 2020) and phase retrieval (Klusowski et al., 2019; Chen et al., 2013)) of mixture models such as MLR and GMM can spur significant interest within the statistics community toward establishing rigorous theoretical foundations for generative diffusion models.
@@ B.3 Proof for Approximate Dynamic Equations
$\alpha^{t}$ $=$ $m(\alpha^{t},\nu^{t})=\mathbb{E}[\tanh(\alpha^{t}X+\nu^{t})X]=\alpha^{t}(1-[\beta^{t}]^{2})+(1-[\beta^{t}]^{2})\mathcal{O}([\alpha^{t}]^{3}),$
$\beta^{t}$ $=$ $n(\alpha^{t},\nu^{t})=\mathbb{E}[\tanh(\alpha^{t}X+\nu^{t})]=\beta^{t}\{1-[\alpha^{t}]^{2}(1-[\beta^{t}]^{2})\}+\beta^{t}(1-[\beta^{t}]^{2})\mathcal{O}([\alpha^{t}]^{4})$
@@ E.1 Modified Log-Sobolev Inequality
$\operatorname{Ent}_{\mu}[f]$ $=$ $\left\{\begin{array}[]{ll}\int f\log f\mathrm{d}\mu-\left(\int f\mathrm{d}\mu\right)\log\left(\int f\mathrm{d}\mu\right)=\int f\log\frac{f}{\int f\mathrm{d}\mu}\mathrm{d}\mu&\operatorname{if}\int f\log(1+f)\mathrm{d}\mu<\infty\\ \infty&\operatorname{otherwise}\end{array}\right$.

### v2

@@ 1.1 Related Work
When ground truth parameters of some components in mixture models have no separation, this means we are using a model with more components to fit a ground truth data distribution with fewer components, which is called the overspecified setting. In particular, the 2GMM/2MLR model is considered overspecified when there is no separation between the true regression parameters or location parameters of its two components. Moreover, it has been observed that the convergence of EM can be prohibitively slow when dealing with poorly separated mixtures (Redner & Walker, 1984), while achieving rapid linear convergence or superlinear convergence in cases with strong separation (Kwon et al., 2021; Ghosh & Kannan, 2020). The convergence rate of regression parameters using EM for 2GMM in the overspecified setting, assuming known mixing weights, is analyzed in Dwivedi et al. (2020b); Dhawan et al. (2023). Kwon et al. (2019; 2021) derived iteration complexity and statistical accuracy for the known-weight case across all SNR regimes—but their results in the low-SNR setting apply only when mixing weights are known to be balanced. Kwon & Caramanis (2020) analyzed EM for MLR with more than two components under restrictive conditions—such as requiring initializations very near the ground truth and well-separated regression parameters—which do not cover cases of overspecification where the ground truth parameters coincide. Nevertheless, there is a lack of prior research formally analyzing the evolution of EM updates for 2MLR with unknown regression parameters and mixing weights in the overspecified setting, with most prior research primarily concentrating on scenarios with known mixing weights. The behavior of EM updates in overspecified settings is not fully understood when dealing with unknown mixing weights, whether they are balanced or unbalanced. In this paper, we address this gap by conducting a detailed analysis of the EM updates for overspecified 2MLR, based on the expressions of integrals with Bessel functions given in (Luo & Hashemi, 2024). Even though Luo & Hashemi (2024) provided closed-form EM update rules (using Bessel functions) for 2MLR across all SNR regimes, but they focus primarily on the high-SNR case and do not discuss convergence guarantees or the dynamics in the overspecified setting and the low-SNR regime. The statistical rates for parameter estimation of a class of overspecified finite mixture models have been studied in several studies (Chen, 1995; Ho & Nguyen, 2016; Heinrich & Kahn, 2018; Doss et al., 2020; Manole & Ho, 2020; 2022; Ho et al., 2022b; Nguyen et al., 2023a). The optimal minimax rates for 2MLR with known balanced mixing weights have been determined in Kwon et al. (2021), while those for 2GMM with known mixing weights have been established in Dwivedi et al. (2020b) using the scheme outlined by Balakrishnan et al. (2017). However, the statistical error for 2MLR with unknown mixing weights remains unclear. We develop new techniques to sharpen the analysis for this statistical error.
@@ 3 Motivating Examples
As motivating examples, we highlight the tasks of haplotype assembly in bioinformatics and genomics (Cai et al., 2016) and phase retrieval, which arises in numerous fields including acoustics, optics, and quantum information (Candes et al., 2015), and also learning overparameterization models and Mixture of Experts (MoE) models, to justify the practical implications of our work.
@@ 3.1 Haplotype Assembly
Haplotypes are sequences of chromosomal variations in an individual’s genome that are crucial for determining the individual’s disease susceptibility. Haplotype assembly involves reconstructing these sequences from a mixture of sequenced chromosome fragments. Notably, humans have two haplotypes, i.e., they are diploid organisms (see Cai et al. (2016) for a clear mathematical formulation of the problem). For diploids, the primary challenge is to reconstruct two distinct haplotypes (binary sequences of single nucleotide polymorphisms—SNPs) from short, noisy sequencing reads. Each read corresponds to a local window of the genome but originates from one of the two chromosomes. The ambiguity in the haplotype origin of each read can be modeled as a mixture of two linear regression models with symmetric parameters, aligning with the model discussed in our work. Following Cai et al. (2016), let $\theta^{\ast}\in\{-1,+1\}^{d}$ represent one haplotype, and the other haplotype is its negative, $\theta^{\ast}$. The binary variable $z_{i}\in\{-1,+1\}$ indicates the haplotype origin of the $i$-th read, where the probability of the read originating from $\theta^{\ast}$ is $\mathbb{P}[z_{i}=+1]=\pi^{\ast}(1)$, and the probability of it originating from $\theta^{\ast}$ is $\mathbb{P}[z_{i}=-1]=\pi^{\ast}(2)$. Furthermore, the $j$-th entry, $\varepsilon_{i}[j]$ for $j\in[d]$, of the noise vector $\varepsilon_{i}$ follows a distribution defined by a fixed error probability $p_{e}$: specifically, the noise causes an error (a flip) with probability $p_{e}$, meaning $\mathbb{P}(\varepsilon_{i}[j]=-2z_{i}\theta^{\ast}[j])=p_{e}$, and the noise is zero (correct reading) with probability $1-p_{e}$, meaning $\mathbb{P}(\varepsilon_{i}[j]=0)=1-p_{e}$. Given this framework, the read signal $y_{i}$ can be modeled by the following two-mixture model: $y_{i}=z_{i}\theta^{\ast}+\varepsilon_{i}$. The primary goal is to estimate the unknown ground truth parameters, which are the mixing probabilities $\pi^{\ast}=(\pi^{\ast}(1),\pi^{\ast}(2))$ and the haplotype $\theta^{\ast}$, using the dataset of read signals $\{y_{i}\}_{i=1}^{n}$. It should be noted that while this is also a two-mixture model, it features a distinct formulation and noise structure.
@@ 3.2 Phase Retrieval
$y^{\prime}_{i}:=y_{i}^{2}-\sigma^{2}=|\langle x_{i},\theta^{\ast}\rangle|^{2}+\xi_{i}.$ (9)
This formulation is essentially the phase retrieval model with a heteroskedastic error term $\xi_{i}:=2(-1)^{z_{i}+1}\langle x_{i},\theta^{\ast}\rangle\varepsilon_{i}+(\varepsilon_{i}^{2}-\sigma^{2})$, which has zero mean, i.e., $\mathbb{E}[\xi_{i}]=0$. Therefore, by leveraging our results on symmetric 2MLR, we can directly establish convergence guarantees for the corresponding phase retrieval problem. For phase retrieval problems, several theoretical guarantees for the parameter estimate $\hat{\theta}$ have been established. Regarding the convex formulation (Chen et al., 2015), for $n\gtrsim d$ samples, the relative error bound is $\|\hat{\theta}-\theta^{\ast}\|/\sigma\lesssim\sqrt{d/n}\log^{4}n+\min(\sqrt{d/n}/\eta,\left(d/n\right)^{1/4})\log^{4}n$, where $\eta=\|\theta^{\ast}\|/\sigma$ represents the signal-to-noise ratio, as shown on page 10 of Chen et al. (2015). In the case of agnostic estimation (Neykov et al., 2016), for $n\gtrsim d^{2}\log d$ samples, the error bound of the estimate $\hat{\theta}$ satisfies $\|\hat{\theta}-\theta^{\ast}\|/\|\theta^{\ast}\|\lesssim\sqrt{(d\log d)/n}$ under the constraint $\|\hat{\theta}\|=\|\theta^{\ast}\|$, as established on pages 3 and 9 of Neykov et al. (2016). For the EM algorithm when the sample size $n$ is sufficiently greater than the dimension $d$ ($n\gtrsim d$), the relative error bound for the estimated vector $\hat{\theta}$ depends on the initialization of the mixing weight $\pi^{0}$. With unbalanced initialization of mixing weights $\|\pi^{0}-\frac{1}{2}\|_{1}\gtrsim(d/n)^{1/4}$, the relative error is bounded by $\|\hat{\theta}-\theta^{\ast}\|/\sigma\lesssim\sqrt{d/n}$. Conversely, with balanced initialization of mixing weights $\|\pi^{0}-\frac{1}{2}\|_{1}\lesssim(d/n)^{1/4}$, the relative error bound is less favorable, given by $\|\hat{\theta}-\theta^{\ast}\|/\sigma\lesssim(d/n)^{1/4}$ (see Theorem 6.1). Regarding sample complexity, our results require $n\gtrsim d$ samples, matching the sample complexity requirement established in Chen et al. (2015) and improving upon the $n\gtrsim d^{2}\log d$ requirement in Neykov et al. (2016). In terms of error rates, with balanced initialization, our rate $(d/n)^{1/4}$ matches the second term in Chen et al. (2015), demonstrating that the EM algorithm achieves a better rate when SNR $\eta$ is sufficiently small. With unbalanced initialization, our rate $\sqrt{d/n}$ matches the leading term $\sqrt{d/n}\log^{4}n$ in Chen et al. (2015) and improves upon the $\sqrt{d\log d/n}$ rate in Neykov et al. (2016) by removing the logarithmic factor when SNR $\eta\asymp 1$. These theoretical comparisons demonstrate that our EM-based approach achieves competitive or improved error rates compared to existing phase retrieval methods, while providing explicit characterization of the initialization-dependent convergence behavior.
3.3 Overparametrization Models
For the general setting of overparameterization models, our theoretical results provide one of the fundamental examples. In our theoretical results of the EM algorithm for the 2MLR model, we exhibit linear convergence with an unbalanced guess of mixing weights, while showing sublinear convergence ($\alpha^{t}\asymp 1/\sqrt{t}$) with a balanced guess of mixing weights (see Theorems 5.16.1), which is consistent with the convergence rate of $\mathcal{O}(1/\sqrt{t})$ for the overparameterized Gaussian mixture model (Xu et al., 2024). Interestingly, for the problem of low-rank matrix factorization (Xiong et al., 2023) in the overparameterization regime, an exponentially faster linear convergence rate is achieved using gradient descent with an asymmetric parameterization, while gradient descent with symmetric parameterization exhibits a sublinear rate of $1/t^{2}$. This suggests that imbalance accelerates the convergence rate. Overparameterization also has an impact on the convergence rate of gradient descent for learning a single neuron in neural networks (Xu & Du, 2023). The method exhibits linear convergence in the exact-parameterization regime, but shows a sublinear convergence rate of $1/t^{3}$ in the overparameterization regime. Therefore, overparameterization can exponentially slow down the convergence rate of Gradient Descent (GD).
3.4 Mixture of Experts
While the EM algorithm has been shown to be a powerful tool for learning Mixture of Experts (MoE) models (Fruytier et al., 2025), establishing the convergence rates of the Maximum Likelihood Estimator (MLE) for these complex mixture models under exact-specified and over-specified settings remains a significant open challenge (Ho et al., 2022a; Nguyen et al., 2023b; 2024). In this context, our theoretical analysis of the EM algorithm for the 2MLR model serves as a fundamental example. By establishing rigorous guarantees in this setting, our work provides the necessary theoretical groundwork to deepen the understanding of convergence behaviors in more complex architectures, such as MoE and deep mixture models.
@@ 5 Population Level Analysis
Intuitive Explanation of Theorem 5.1. For GMM/MLR models, running the EM algorithm is equivalent to performing a step of Gradient Descent on the negative log-likelihood (Kwon et al., 2024). When an unbalanced estimate of mixing weights is given, the negative log-likelihood retains its dominant quadratic term, $\alpha]^{2}$. This quadratic form ensures that the EM algorithm behaves like a Gradient Descent step on a strongly convex function, suggesting a fast, linear convergence rate for $\alpha^{t}$. In contrast, when a balanced estimate of mixing weights is given, the $\alpha]^{2}$ term in the negative population log-likelihood cancels out (page 22 of Luo & Hashemi (2024) and Lemma A.3). Specifically, the likelihood term can be expanded as $\alpha]^{2}/2-\mathbb{E}[\ln\cosh(\alpha X)]=[\alpha]^{2}/2-\mathbb{E}[(\alpha X)^{2}/2-(\alpha X)^{4}/12+\ldots]\approx 3[\alpha]^{4}/4$. The EM algorithm, being equivalent to Gradient Descent, then follows a path approximated by $\alpha^{t+1}\approx\alpha^{t}-\nabla(3[\alpha^{t}]^{4}/4)=\alpha^{t}-3[\alpha^{t}]^{3}$, leading to a much slower, sublinear convergence rate where $\alpha^{t}$ is proportional to $1/\sqrt{t}$.
@@ 6 Finite-Sample Level Analysis
Intuitive Explanation of Theorem 6.1. The final statistical accuracy of the MLE estimate is governed by the invertibility of the Fisher Information matrix, which is the second-order derivative of the negative population log-likelihood. With an unbalanced estimate of mixing weights, the negative log-likelihood maintains the $\alpha]^{2}=\|\theta\|^{2}/\sigma^{2}$ term, which ensures the Fisher Information matrix (w.r.t. the regression parameters) is invertible. As a well-established result (Van der Vaart, 2000), this invertibility guarantees that the MLE estimate achieves the standard parametric rate of order $n^{-1/2}$, and therefore the final accuracy of the EM algorithm is also proportional to $n^{-1/2}$. However, when a balanced estimate is given, the critical term $\alpha]^{2}$ vanishes, causing the Fisher Information matrix to be singular. This singularity slows down the standard parametric rate of final accuracy (Dwivedi et al., 2020b), resulting in a converged $\alpha^{T}$ that exhibits a rate proportional to $n^{-1/4}$.
@@
Proposition 6.4 (see proof in Appendix F, Subsection F.4) demonstrates our sample complexity $n=\Omega(d\vee\log\frac{1}{\delta})$, whereas Lemma 1 of Dwivedi et al. (2020b) presents a sample complexity $n=\Omega(d\log\frac{1}{\delta})$. The difference lies in the techniques used: their multiplicative log factor is derived through the standard symmetrization technique with Rademacher variables and the application of the Ledoux-Talagrand contraction inequality (see the proof of Lemma 1 on page 44 of Dwivedi et al. (2020b) and Appendix E of Kwon et al. (2021) for Lemma 11). In contrast, we obtain our additive log factor by applying the rotational invariance of Gaussians and expressing the $\ell_{2}$ norm of the orthogonal error as the geometric mean of two Chi-square distributions (see Appendix E, Subsection E.3).
@@ 8 Experiments
In this section, we validate our theoretical findings in the previous sections with numerical experiments. The code for our numerical experiments is available at https://github.com/dassein/em_overspecified_mlr.
@@ 9 Conclusions
Building upon the analysis and established connections between the diffusion model objective and the classic EM algorithm in GMM (Shah et al., 2023), we foresee extending this analysis from GMM to MLR and establishing the time and sample complexities involved in learning the diffusion model objective, as discussed in recent works (Chen et al., 2024a; Gatmiry et al., 2024). The practical applications (haplotype assembly (Cai et al., 2016; Sankararaman et al., 2020) and phase retrieval (Klusowski et al., 2019; Chen et al., 2015)) of mixture models such as MLR and GMM can spur significant interest within the statistics community toward establishing rigorous theoretical foundations for generative diffusion models.
@@ B.3 Proof for Approximate Dynamic Equations
$\alpha^{t+1}$ $=$ $m(\alpha^{t},\nu^{t})=\mathbb{E}[\tanh(\alpha^{t}X+\nu^{t})X]=\alpha^{t}(1-[\beta^{t}]^{2})+(1-[\beta^{t}]^{2})\mathcal{O}([\alpha^{t}]^{3}),$
$\beta^{t+1}$ $=$ $n(\alpha^{t},\nu^{t})=\mathbb{E}[\tanh(\alpha^{t}X+\nu^{t})]=\beta^{t}\{1-[\alpha^{t}]^{2}(1-[\beta^{t}]^{2})\}+\beta^{t}(1-[\beta^{t}]^{2})\mathcal{O}([\alpha^{t}]^{4})$
@@ E.1 Modified Log-Sobolev Inequality
$\operatorname{Ent}_{\mu}[f]$ $=$ $\left\{\begin{array}[]{ll}\int f\log f\mathrm{d}\mu-\left(\int f\mathrm{d}\mu\right)\log\left(\int f\mathrm{d}\mu\right)=\int f\log\frac{f}{\int f\mathrm{d}\mu}\mathrm{d}\mu\quad&\operatorname{if}\int f\log(1+f)\mathrm{d}\mu<\infty\\ \infty&\operatorname{otherwise}\end{array}\right$.